# Signs as Tokens: An Autoregressive Multilingual Sign Language Generator

### v1

@@ 4.3 User Study
The ultimate goal of sign language generation research is to bridge the communication gap between deaf and hearing individuals. In addition to the objective metrics presented in the previous tables, we conducted a user study involving five professional ASL signers. These signers were asked to rate the alignment between generated signs and text annotations on a scale from 1 to 10, with higher scores indicating better semantic conveyance of the texts. Ground truth sign motions were also included in the study. Specifically, we provided 15 generated signs from both the baseline, MotionGPT, and our proposed method, SOKE. The order of the signs was randomly shuffled to prevent potential bias.
As shown in Figure 7, our method achieved an encouraging average rating of 5.24, significantly surpassing the baseline, which received a low score of 1.36. These results align with the above qualitative and quantitative evaluations, indicating the potential of our method in constituting a two-way communication system between the deaf and hearing.
Figure 7: User study with professional ASL signers. We report average ratings of the baseline (MotionGPT), our proposed SOKE, and ground truth motions.
@@ 5 Conclusion
In this paper, we introduce SOKE, an autoregressive multilingual sign language generator. Unlike existing approaches that treat sign language generation as a visual content generation task, we focus on the linguistic nature of sign languages, characterized by discrete semantic units. To this end, we develop a specialized sign motion tokenizer that maps continuous motions to discretized tokens in a latent space. Additionally, inspired by the multi-channel characteristics of sign languages, we decompose the tokenizer into three parallel components to independently model different body parts. Once sign motions are discretized, we utilize a pretrained multilingual LM to perform autoregressive generation using a set of customized special tokens. Furthermore, we curate CSL-Daily with 3D pose annotations to support multilingual SLG. Extensive qualitative and quantitative evaluations on a combined ASL-CSL dataset demonstrate the effectiveness of our proposed method.
Acknowledgements.
S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1), EPSRC Project GNOMON (EP/X011364/1) and Turing AI Fellowship (EP/Z534699/1). R.A. Potamias was supported by EPSRC Project GNOMON (EP/X011364/1). We would also like to thank anonymous signers who participated in the user study.
Supplementary Material
Appendix A Curating CSL-Daily with SMPL-X Pose
Figure S1: Qualitative comparisons between ground truth video frames and reconstructed meshes obtained from the proposed SMPL-X pose fitting pipeline on CSL-Daily.
@@ Appendix A Curating CSL-Daily with SMPL-X Pose
Since CSL-Daily [79] does not provide 3D annotations, we perform qualitative evaluations, as illustrated in Figure S1. The results clearly demonstrate that the proposed pose fitting pipeline can accurately reconstruct 3D hands and is robust across various handshapes. To quantitatively assess the pipeline, we further apply it to the SGNify mocap dataset [11], which includes 57 signs with annotated meshes. The results presented in Table S1 indicate that our method achieves the lowest hand reconstruction errors and comparable body errors to the previous best method [3], establishing our approach as a powerful tool for curating more sign language datasets in the future.
Appendix B More Qualitative Results
Please refer to the enclosed .mp4 file for video demonstrations of generated signs. These demos include ground truth sign videos, as well as generations from the baseline method, MotionGPT [22], and our proposed SOKE. Additionally, we present several qualitative results showcasing both generated signs (Figure S2) and reconstructed signs (Figure S3).
(a) CSL-Daily.
(b) How2Sign.
Figure S2: Qualitative comparisons of generated signs between our proposed method, SOKE, with the baseline method, MotionGPT [22], on CSL-Daily and How2Sign. The results indicate that SOKE can generate finer-grained hand articulations than MotionGPT.
Figure S3: Qualitative comparisons of reconstructed signs on CSL-Daily (first four columns) and How2Sign (last four columns). The results verify that using our proposed decoupled tokenizer (DETO) significantly enhances the reconstruction quality of hands.
Appendix C More Quantitative Results
$N_{Z}^{B}$ $N_{Z}^{LH}=N_{Z}^{RH}$ How2Sign CSL-Daily
@@ Appendix C More Quantitative Results
All Hand Face Body Hand All Hand Face Body Hand
96 128 23.35 6.73 1.13 19.35 6.67 28.24 5.25 1.18 23.46 5.21
96 256 23.37 6.52 1.11 19.32 6.70 28.33 5.13 1.19 23.32 5.36
64 192 24.28 6.53 1.14 19.72 6.71 29.52 5.13 1.20 24.42 5.23
128 192 23.97 6.55 1.11 19.69 6.81 28.67 5.16 1.23 23.51 5.31
96 192 23.26 6.47 1.11 19.00 6.65 28.10 5.10 1.17 23.27 5.20
Table S2: Study on the codebook sizes for the body ($N_{Z}^{B}$) and hands ($N_{Z}^{LH},\ N_{Z}^{RH}$).
Codebook Size. We perform a hyper-parameter analysis on the codebook sizes for the body ($N_{Z}^{B}$) and hands ($N_{Z}^{LH},\ N_{Z}^{RH}$). As shown in Table S2, we find that using either larger or smaller codebooks results in degraded reconstruction performance. Our default configuration ($N_{Z}^{B}=96$, $N_{Z}^{LH}=N_{Z}^{RH}=192$) delivers the best performance among all settings.
@@
Body L-Hand R-Hand Body L-Hand R-Hand
<$\mathtt{bos}$> 12.03 6.19 5.94 12.55 6.08 6.37
<$\mathtt{Lang\_p}$>(default) 7.61 2.83 2.93 8.08 2.20 2.66
Table S3: Study on the start tokens for part-wise decoding.
Part-wise Decoding. To integrate the language model (LM) with the decoupled tokenizer, we develop a part-wise decoding strategy where each decoding step begins with a special token indicating the body part. To verify its effectiveness, we conduct an experiment by replacing the start token with <$\mathtt{bos}$>. As shown in Table S3, using <$\mathtt{bos}$>as the start token fails to provide the LM with body part information, leading to significant performance degradation.
Appendix D Broader Impacts and Limitations
@@
We have validated the proposed multilingual sign language generator on two widely-adopted sign languages: Chinese Sign Language, which is used by approximately 4 million people in China, and American Sign Language, with around half a million users in North America [63]. In the future, we plan to extend our method to support additional sign languages, such as British Sign Language [1] and Indian Sign Language [25].

### v3

@@ 4.3 User Study
In addition to the objective metrics presented in the previous tables, we conducted a user study involving 5 professional ASL signers and 4 professional CSL signers. These signers were asked to rate the alignment between generated signs and text annotations on a scale from 1 to 10, with higher scores indicating better semantic conveyance of the texts. Ground truth sign motions were also included in the study. Specifically, we provided 15 generated signs from both the baseline, S-MotionGPT, and our proposed method, SOKE. The order of the signs was randomly shuffled to prevent potential bias.
As shown in Figure 7, our method achieved an encouraging average rating of 5.32/6.45, significantly surpassing the baseline, which received a low score of 1.36/2.70 on How2Sign and CSL-Daily, respectively. These results align with the above qualitative and quantitative evaluations, indicating the potential of our method in constituting a two-way communication system between the deaf and hearing.
Figure 7: User study with professional ASL and CSL signers. We report average ratings of S-MotionGPT, our proposed SOKE, and ground truth motions.
@@ 5 Conclusion
In this paper, we propose SOKE, a retrieval-enhanced multilingual sign language generator. Unlike existing methods that view sign language generation as a visual content generation task, we emphasize its linguistic nature, characterized by discrete semantic units. To capture the multi-cue characteristics of sign languages, we employ a decoupled tokenizer to map continuous motions into discrete tokens for the upper body and both hands. To balance efficiency and generation quality, we propose a novel multi-head decoding method that reduces decoding steps by two-thirds while dynamically integrating information from different body parts. Additionally, we present a simple yet effective retrieval-enhanced approach by leveraging external sign dictionaries. Extensive evaluations on a combined ASL-CSL-DGS dataset validate the effectiveness of our method.
Acknowledgements. S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1), EPSRC Project GNOMON (EP/X011364/1) and Turing AI Fellowship (EP/Z534699/1). R.A. Potamias and R. Zuo were supported by EPSRC Project GNOMON (EP/X011364/1). This research was also supported by the Sovereign AI AIRR programme. We would also like to thank anonymous signers who participated in the user study.
Appendix A Curating SMPL-X Poses
@@
Figure S9: Qualitative comparisons of generated signs between our proposed method, SOKE, with the SOTA method, S-MotionGPT [27], on the test sets of How2Sign (left), CSL-Daily (middle), and Phoenix-2014T (right).
Since neither CSL-Daily [89] nor Phoenix-2014T [4] provides 3D annotations, we perform qualitative evaluations, as illustrated in Figure S8. The results clearly demonstrate that the proposed pose fitting pipeline can accurately reconstruct 3D hands and is robust across various handshapes. To quantitatively assess the pipeline, we further apply it to the SGNify mocap dataset [16], which includes 57 signs with annotated meshes. The results presented in Table S5 indicate that our method achieves the lowest hand reconstruction errors and comparable body errors to the previous best method [3], establishing our approach as a powerful tool for curating more sign language datasets in the future.
Appendix B Additional Qualitative Results
Please refer to our project page for video demonstrations of generated signs. These demos include ground truth sign videos, as well as generations from the SOTA method, S-MotionGPT [27], and our proposed SOKE. Additionally, we provide several qualitative results to showcase the generated signs (Figure S9) and highlight the effectiveness of our retrieval-enhanced SLG approach (Figure S10).
Figure S10: Qualitative ablation study for retrieval-enhanced SLG. (Left: How2Sign; Middle: CSL-Daily; Right: Phoenix-2014T.)
Appendix C Additional Quantitative Results
@@
Body Hand Body Hand Body Hand
96 128 19.37 7.07 23.52 5.80 25.79 7.35
96 256 19.37 6.86 23.52 5.52 25.79 7.11
64 192 20.04 6.65 24.13 5.13 26.02 6.78
128 192 19.95 6.65 23.91 5.13 26.27 6.78
96 192 19.37 6.65 23.52 5.13 25.79 6.78
Table S6: Study on the codebook sizes for the body ($N_{Z}^{B}$) and hands ($N_{Z}^{LH},\ N_{Z}^{RH}$). We use procrustes-aligned mean per joint position error (PA-MPJPE) to assess the reconstruction performance of the decoupled tokenizer.
@@
Body Hand Body Hand Body Hand
0.1 7.95 2.82 7.46 2.13 5.47 2.04
0.2 7.28 2.76 6.91 1.95 5.08 1.68
1/3 6.82 2.35 6.24 1.71 4.77 1.38
0.4 7.34 2.62 7.11 1.91 6.39 1.96
Table S7: Study on the impact of $\lambda$, a hyper-parameter used for fusing part-wise token embeddings in our multi-head decoding method.
Codebook Size. We perform a hyper-parameter analysis on the codebook sizes for the body ($N_{Z}^{B}$) and hands ($N_{Z}^{LH},\ N_{Z}^{RH}$) in our decoupled tokenizer. As shown in Table S6, we find that using either larger or smaller codebooks results in degraded reconstruction performance. Our default configuration ($N_{Z}^{B}=96$, $N_{Z}^{LH}=N_{Z}^{RH}=192$) delivers the best performance among all settings.
Impact of $\lambda$ on SLG. In our multi-head decoding method, we introduce a hyper-parameter, $\lambda$, to control the weight of hand tokens during embedding fusion. The results in Table S7 demonstrate that $\lambda=1/3$, i.e., assigning equal weights to the body and hands, yields the best performance. This further underscores the importance of each body part in conveying the semantics of sign languages.
[…]
Avg Body Hand Avg Body Hand Avg Body Hand
S-MotionGPT × 5.91 11.23 4.39 5.34 10.81 3.78 4.75 9.45 3.41
Ours × 4.14 7.92 3.07 4.18 8.18 3.04 3.83 7.25 2.85
Ours 3.34 6.82 2.35 2.72 6.24 1.71 2.13 4.77 1.38
Table S8: Performance of our method on monolingual datasets.
Monolingual Performance. As shown in Table S8, our method still outperforms the SOTA method, S-MotionGPT, when training on monolingual SL datasets, while the best results are achieved by the multilingual version of our method.
Appendix D Illustration of Decoupled Tokenizer
Figure S11: Workflow of our decoupled tokenizer. It is composed of three parallel VQ-VAEs, each dedicated to generating motion tokens for a different part of the signer’s body: the upper body, left hand, and right hand.
As shown in Figure S11, we provide an illustration of our decoupled tokenizer for better understanding. It utilizes three VQ-VAEs to model the key regions of a signer: the upper body, left hand, and right hand.
Appendix E Discussion
@@
Future Works. We have validated the proposed multilingual sign language generator on three widely-adopted sign languages, Chinese, American, and German sign language [89, 11, 34]. As the scalability of our approach has been demonstrated in Table 3 of the main paper, in the future, we plan to extend our method to support more sign languages, such as British Sign Language [1] and Indian Sign Language [31].